OpenStudy (rea201):

Help Calculus How many ways can 9 numbers be arranged on a small rotating wheel relative to a fixed point? Consider a basket that contains 15 slips of paper numbered from 1 to 15. Two slips of paper are drawn at random. What is the probability of drawing 2 even numbers?

OpenStudy (kenljw):

9 8 7 ... 1 = 9! 1,2,...,15 15 numbers, n odd,n-1even n + n-1=15, n=8 first draw P(E)=7/15. P(O)=8/15 Conditional Probabilities second draw P(E|E)=6/14, P(E|O)=7/14. P(O|E)=8/14, P(O|O)=7/14 it's been awhile since I've done these, try Baye's theorem

OpenStudy (rea201):

thanks!!!!